.. 10.22 A hotel housekeeper, Alison Harvey, was observed five times on each of four task elements, as shown in the followingPlease show a step by step answer to this problem with all possible calculations and show pictures if any tables/graphs were used. Thanks!
Answer #1

In this question no table and graph is applicable a detailed explanation is given below as:

(i)For Check minibar: The Observed mean time=1.5 minutes(1.5+1.6+1.4+1.5+1.5/5)

Basic time is formulated as by the formula as (Observed time X Performance Rating)/100

By calculation of Basic time=100 x 1.5/100=1.5 minutes

Standard Time = Basic Time + Allowances

As allowance is 10%

Standard time=1.1[1.5]=1.65 minutes for Check minibar.

(ii)Make one bed: Observed mean time=2.3 minutes(2.3+2.5+2.1+2.2+2.4/5)

Basic time is formulated as by the formula as (Observed time X Performance Rating)/100

By calculation of Basic time=90 x 2.3/100=2.07 minutes

Standard Time = Basic Time + Allowances

As allowance is 10%

Standard time=1.1[2.07]=2.277 minutes to make one bed.

(iii)Vacuum floor: Observed mean time=1.7 minutes(1.7+1.9+1.9+1.4+1.6/5)

Basic time is formulated as by the formula as (Observed time X Performance Rating)/100

By calculation of Basic time=120 x 1.7/100=2.04 minutes

Standard Time = Basic Time + Allowances

As allowance is 10%

Standard time=1.1[2.04]=2.244 minutes for Vacuum floor.

(iv)Clean bath: Observed mean time=3.5 minutes(3.5+3.6+3.6+3.6+3.2/5)

Basic time is formulated as by the formula as (Observed time X Performance Rating)/100

By calculation of Basic time=100 x 3.5/100=3.5 minutes

Standard Time = Basic Time + Allowances

As allowance is 10%

Standard time=1.1[3.5]=3.85 minutes for Clean bath.

So the standard time for whole process is (i)+(ii)+(iii)+(iv)=10.021 minutes